OpenStudy (anonymous):

The sum of two number is -11. Twice the first number minus the second is 32 Find the numbers. System of Equations Applications! Please help

Directrix (directrix):

Let x be one number and y the other 1) x + y = -11 2) 2*x - y = 32

Directrix (directrix):

@28Tylerr Add the two equations. The y variables should sum to zero. Then, you can solve for x.

OpenStudy (anonymous):

Can you explain that into detail steps?

Directrix (directrix):

Explain what? How I got the equations or how to solve them?

OpenStudy (anonymous):

How to solve them

Directrix (directrix):

1) x + y = -11 2) 2*x - y = 32 I'm going to write it out by hand. Hold on.

Directrix (directrix):